In testing whether the means of two normal populations are equal, summary statistics computed for two independent samples are as follows:

Brand X Brand Y

๐‘›1 = 25 ๐‘›2 = 25

x=7.3 x=6.8

S1=1.05 S2=1.20

Assume that the population variances are equal. Then the 95% confidence interval for the difference between population means of band ๐‘‹ and brand ๐‘Œ is:

1. (-0.1407; 1.1407)

2. (0.5; 0.0906)

3. (0.490; 0.9506)

4. (0.3976; 0.5609)

5. (0.5906; 0.4094)


1
Expert's answer
2021-09-21T10:47:25-0400

Solution:

"𝑛_1 = 25, 𝑛_2 = 25
\\\bar x_1=7.3, \bar x_2=6.8
\\S_1=1.05, S_2=1.20"

"H_0:\mu_1= \mu_2
\\ H_1:\mu_1\ne \mu_2"

ฮฑ = 0.05ย ย ย ย ย ย 

Degrees of freedom = 25 + 25 - 2 = 48

Lower Critical t- score = -2.011ย ย ย ย ย 

Upper Critical t- score = 2.011

Now, we reject "H_0" if "|t|>2.011"

Pooled variance, "S_p=\sqrt{\dfrac{(n_1-1)S_1^2+(n_2-1)S_2^2}{n_1+n_2-2}}"

"S_p=\sqrt{\dfrac{(24)(1.05)^2+(24)(1.2)^2}{48}}=1.127"

Then, SE"=S_p\sqrt{\dfrac{1}{n_1}+\dfrac{1}{n_2}}=1.127\sqrt{\dfrac{2}{25}}=0.318"

Now, 95% CI"=(\bar x_1-\bar x_2)\pm t_{0.05,48}\ SE"

"=(7.3-6.8)\pm2.011(0.318)
\\=(-0.139,1.139)
\\\approx(-0.14, 1.14)"

Thus, option 1 is correct.